# AAPM: Large Language Model Agent-based Asset Pricing Models

### v1

@@ 4 Experiment
We conduct experiments to assess the asset pricing efficacy of the proposed AAPM. The experimental setup is detailed in Section 4.1. Subsequently, we present the outcomes of the portfolio optimization experiments in Section 4.2 and the asset pricing error in Section 4.3. An extensive ablation study of our method is provided in Section 4.4. Furthermore, we explore the predictive capabilities of refined news on economic indicators and stock movements in Appendix C.
@@ 4.1 Experiment Setting
We build a dataset comprising two years of WSJ articles spanning from September 29, 2021, to September 29, 2023, following the knowledge cutoff of the version of GPT we used. This approach mitigates potential information leaks while maintaining continuity in note $n$. Besides the LLM filtering described in Section 3.1, we also manually excluded articles on unrelated topics like travel, lifestyle, and puzzles, based on their WSJ categories. Visualizations of our news dataset can be found in Appendix B. The daily asset returns are sourced from CRSP, while daily risk-free returns and market returns are obtained from Kenneth French’s data library  [^2] https://mba.tuck.dartmouth.edu/pages/faculty/ken.french.
@@
We select five recent asset pricing baselines from highly reputed financial economics journals, validated under current empirical finance standards, as indicated by Jensen et al. (2023), to assess our approach: NN Gu et al. (2020) introduced a deep neural network for asset pricing; IPCA Kelly et al. (2020) developed an instrumental PCA to identify hidden factors and loadings; CA Gu et al. (2021) proposed to use a conditional autoencoder; NF Bybee et al. (2023) employs LDA for the WSJ news as hidden factors similar to ours; and CPZ Chen et al. (2024) utilized GAN to address stochastic discount factors. We replicated these models using the configurations from their respective papers with their carefully chosen factor sets. For both our method and the baselines, we performed a hyper-parameter search to compare the best results. The hyper-parameter optimization setting for our method is detailed in Appendix A.
@@ 4.2 Portfolio Optimization
NN 3.82 2.83 2.36 4.82 8.12 9.12
IPCA 4.07 2.96 2.66 3.77 5.77 8.63
CA 4.03 2.85 2.55 3.79 6.31 4.66
NF 3.73 2.76 2.34 5.12 7.91 6.31
CPZ 4.10 3.02 2.61 4.32 6.27 5.71
Ours 4.38 3.29 3.01 3.66 5.64 5.17
w/ G.4 4.45 3.43 3.09 3.82 5.57 4.77
Table 1: Sharpe Ratio (SR) and Maximal Drawdown (MDD) for Tengency Portfolio (TP), Equal-Weighted (EW) and Value-Weighted (VW) long-short portfolio built based on NN Gu et al. (2020), IPCA Kelly et al. (2020), CA Gu et al. (2021), NF Bybee et al. (2023), CPZ Chen et al. (2024), and our method with the default GPT-3.5 or GPT-4. We bolded the best results and underlined the second bests.
We begin by testing the Sharpe ratio for portfolios built on the predicted returns of individual assets. The Sharpe ratio (SR) Sharpe (1998) quantifies the risk-adjusted performance of a portfolio as $S_{p}=\frac{\bar{r}_{p}-\bar{r}_{f}}{\sigma(r_{p})}$, where $r_{f}$ stands for the risk-free return, $r_{p}$ represents the portfolio return and $\sigma$ indicates the standard deviation. Furthermore, we evaluate the maximum drawdown, which is the largest decrease in the total value of the portfolio up to time $T$, expressed as $MDD(T)=\max_{\tau\in(0,T)}[\max_{t\in(0,\tau)}X(t)-X(\tau)]$. Here, $X(\tau)$ is the highest value and $X(t)$ is the lowest value of the portfolio within the time interval $(0,\tau)$.
@@
The experiment results are presented in Table 1. Our approach achieved the highest SR across all three portfolios, with SR improvements of 6.8%, 8.9%, and 13.2% respectively over the best baseline methods (CPZ for TP and EW, IPCA for VW), averaging a 9.6% increase. Additionally, it secured the best or second-best MDD in TP and EW compared to the leading baseline IPCA, with gains of 2.9% and 2.3% respectively. In VW, the MDD underperforms the top baseline CA by 10.9%. However, substituting GPT-3.5 in our model with GPT-4-0613 which has the same knowledge cutoff, resulted in SR improvements of 8.5%, 13.6%, and 16.2% across the three portfolios, and improved MDD levels to gains of 1.3%, 3.5%, and -2.4% relative to the best baselines.
@@ 4.3 Asset Pricing Error
avg $|\alpha|$ avg $|t(\alpha)|$ $\frac{\#|t(\alpha)|>1.96}{\#test\ assets}$ GRS
NN 0.83 2.89 0.64 6.89
IPCA 0.76 2.45 0.55 6.38
CA 0.77 2.63 0.52 6.42
NF 0.89 2.77 0.62 7.32
CPZ 0.74 2.44 0.49 6.77
Ours 0.66 2.40 0.46 6.34
w/ G.4 0.64 2.36 0.46 6.28
Table 2: Asset pricing errors for anomaly portfolios with NN Gu et al. (2020), IPCA Kelly et al. (2020), CA Gu et al. (2021), NF Bybee et al. (2023), CPZ Chen et al. (2024), and our method with GPT-3.5 and GPT-4. We bolded the best results and underlined the second bests.
@@
Table 2 displays the results. Our method secured either the top or second-best performance among all benchmarks. It demonstrates a 10.8% and 13.5% reduction in average $|\alpha|$ for GPT-3.5 and GPT-4 respectively when compared to CPZ, the leading benchmark, along with a 1.6% and 3.3% increase in t-value. Additionally, there is a 6.1% reduction in the proportion of pricing results with a t-value exceeding 1.96 compared to CPZ for both GPT-3.5 and GPT-4, as well as a 0.6% and 1.6% enhancement in the GRS test compared to IPCA.
Figure 4: Cumulative excess return for decile portfolios.
We move forward to assess the proposed method by applying it to the pricing of decile portfolios. This process includes sorting the assets according to their predicted returns and then forming portfolios for each decile. Figure 4 shows the cumulative excess return over time. The figure clearly demonstrates that each decile creates a distinct ranking of returns in the right position, suggesting that the proposed approach effectively predicts returns at various levels.
@@ 4.4 Ablation Study
SR MDD avg $|\alpha|$ avg $|t(\alpha)|$
NF 2.76 7.91 0.89 2.77
+ Factors 2.66 8.82 0.97 2.86
Naive 2.82 6.03 0.88 2.72
+ RAG 2.94 5.89 0.83 2.66
+ Emb. 2.88 6.42 0.86 2.71
Memory 2.99 6.99 0.81 2.64
+ Factors 3.03 7.12 0.79 2.62
Hybrid 3.14 5.59 0.73 2.49
+ Refine 3.26 6.31 0.70 2.46
+ Notes 3.18 6.91 0.74 2.55
Ours 3.29 5.64 0.66 2.40
Table 3: Ablation study of AAPM and comparison with NF Bybee et al. (2023). “Naive” directly produce the analysis report given news and only daily embeddings are inputted to the pricing network. “+ RAG” introduces the external memory and retrieves Top-K items when performing analysis. “+ Emb.” introduces the asset embeddings. “Memory” baseline incorporate both “+ RAG” and “+ Emb.” “+ Factors” introduces the daily manual factors into the pricing network in “Memory”. “Hybrid” baseline pretrained the pricing network of “Memory”. “+ Refine” refines the analysis report iteratively in NN rounds for “Hybrid”. “+ Notes” introduces the macro economics and micro market notes. “Ours” is our method that combines “+ Refine” and “+ Notes” in “Hybrid”.
@@ 4.4.1 Agent Architecture Design
We analyze our architecture in a reverse manner, beginning with a "Naive" agent that generates the analysis report directly from the refined news without any supplementary information or iterative analysis, while the pricing network solely uses the daily embeddings as input. We then incrementally add components to develop stronger baselines until arriving at our method. The results are shown in Table 3, and the baseline illustrations are provided in Appendix D.
@@
Owing to the analytical capabilities and feature extraction proficiency of LLMs, the “Naive” baseline enhances the SR by 2.2% with comparable pricing errors to NF. Incorporating external memory further boosts the SR by 4.3% and decreases the average $|\alpha|$ by 5.7% over “Naive”, highlighting the significance of additional contextual information when interpreting business news. Moreover, asset embedding contributes to a 2.1% increase in SR and a 2.3% reduction in average $|\alpha|$ by introducing asset-specific loadings.
By combining both, the “Memory” baseline enhances the SR of “Naive” by 6.0% and decreases the average $|\alpha|$ by 8.0% with a lower $t$ value. Incorporating the manual factors, the SR saw a slight increase of 1.3%, while the average $|\alpha|$ decreased by 2.5%. In comparison, the performance of NF declined after the introduction of manual factors, which is consistent with the findings of Bybee et al. (2023), where the inclusion of Fama-French factors reduces the SR, which may due to suboptimal interactions between factors and news features.
After pretraining the pricing network with historical factor data, the performance of the “Hybrid” baseline saw a notable enhancement of 5.0% in SR and a 9.9% reduction in the average $|\alpha|$ when compared to the “Memory” baseline. This demonstrates the synergy between manual factors and LLM-generated reports, resulting in a successful non-linear interaction. The improvements from our iterative refinement and long-term notes over the “Hybrid” baseline are 3.8% and 1.3% in SR, 4.1% and a slight negative -1.4% in average $|\alpha|$, respectively with a lower $t$ value and a similar level of MDD. These enhancements collectively yield 4.8% and 9.6% gains in SR and average $|\alpha|$ respectively in our full method compared to a “Hybrid” baseline, underscoring the effectiveness of our agent architecture design.
@@ 4.4.2 Analysis Depth and Width
Figure 5: The Sharpe ratio of equal-weighted portfolios given different number of $K$ and $N$.
We further investigate the depth of the analysis, which is controlled by the number of iterations $N$ to refine the analysis report, and the width, which is determined by $K$, the amount of relevant information to check. The results are shown in Figure 5. We keep one variable constant and test the other. We observe that the agent benefits from more rounds of analysis and a broader range of relevant information overall with a sharp decline in marginal gain after a certain point around $K\times N=15$, likely due to the sufficiency of the provided information. Thus, we test an extreme case where $N=1$ and $K=15$, resulting in the SR dropping to 3.12. This indicates that iterative refinement is necessary, as items retrieved in different rounds of refinement provide diverse information as the query evolves. In contrast, a single retrieval leads to items falling into the same topic, with the value of additional items decreasing rapidly and potentially introducing noise.
@@ 5 Discussion
Our proposed approach presents a promising method to fuse qualitative discretionary investment with quantitative factor-based strategies through the use of LLM agents. Nonetheless, there is still much to investigate regarding additional capabilities of LLM agents that could further enhance asset pricing power. Firstly, internet access and a broader range of information sources, including those available in SocioDojo, may enable the agent to generate more in-depth analyses, as discretionary investment relies on information beyond just news or domain knowledge. Secondly, employing specialized financial LLMs like FinGPT Yang et al. (2023) could further improve the agent’s financial analytical capabilities. Finally, it is crucial to consider multimodal information, such as diagrams and figures, which are frequently presented in financial documents.
@@ 6 Conclusion
In this research, we introduced AAPM, a model that combines qualitative analysis from the LLM agent with quantitative factors in asset pricing. AAPM surpassed established asset pricing methods in multiple evaluations, including portfolio optimization and asset pricing error. Additionally, we performed an in-depth analysis of each component in our agent design. We believe that our study can improve the comprehension of the interaction between discretionary investment and quantitative factor-based models, toward a society with increased economic efficiency.
Limitations
Our experiments only focus on the US market and English news, which may potentially impact model performance in lower-resources languages. In order to exclude the information leak, we can only apply news data after September 2021 which restricts our study to a 2 years period after this time, however, we use a large test split where half of the dataset was applied as the test set to best evaluate how well the proposed method can be generalized beyond the training period. Finally, public information in the stock market includes not only news, but also reports, reports from social networks, academic journals, opinions from experts, etc.; we do not cover these information nor consider multimodal inputs as discussed in Section 5.

### v2

@@ 4 Experiment
We conduct experiments to assess the asset pricing efficacy of the proposed method. The experimental setup is detailed in Section 4.1. Subsequently, we present the outcomes of the portfolio optimization experiments in Section 4.2 and the asset pricing error in Section 4.3. An extensive ablation study of our method is provided in Section 4.4. Furthermore, we explore the predictive capabilities of refined news on economic indicators and stock movements in Appendix D.
@@ 4.1 Experiment Setting
We build a dataset comprising three years of WSJ articles spanning from September 29, 2021, to September 29, 2024, based on the SocioDojo corpus Cheng & Chin (2024). This approach mitigates potential information leaks while maintaining continuity in note $n$. Besides the LLM filtering described in Section 3.1, we also manually excluded articles on unrelated topics like travel, lifestyle, and puzzles, based on their WSJ categories. Visualizations of our news dataset can be found in Appendix C. The daily asset returns are sourced from CRSP, while daily risk-free returns and market returns are obtained from Kenneth French’s data library  [^2] https://mba.tuck.dartmouth.edu/pages/faculty/ken.french.
@@
We apply OpenAI GPT-4o-0806 Achiam et al. (2023) as default LLM, and the latest reasoning model O1-Preview in addition. Following Cheng & Chin (2024), we set a low temprature=0.2 in LLMs for all of our experiments, balancing deterministic and flexibility in analysis. We select five recent asset pricing baselines from highly reputed financial economics journals, validated under current empirical finance standards, as indicated by Jensen et al. (2023), to assess our approach: Gu et al. (2020) introduced a deep Neural Network for Asset Pricing (NNAP); Kelly et al. (2020) developed an Instrumental PCA (IPCA) to identify hidden factors and loadings; Gu et al. (2021) proposed to use a Conditional Autoencoder (CA); Bybee et al. (2023) employs LDA for the WSJ news as hidden News Factors (NF) similar to ours; and Chen et al. (2024) utilized GAN to address Stochastic Discount Factors (SDF-GAN). We replicated these models using the configurations from their respective papers with their carefully chosen factor sets. For both our method and the baselines, we performed a hyper-parameter search to compare the best results. The hyper-parameter optimization setting for our method is detailed in Appendix A.
@@ 4.2 Portfolio Optimization
NNAP 3.59 2.55 2.29 4.89 7.72 8.99
IPCA 3.85 2.76 2.48 4.42 6.32 8.24
CA 3.77 2.69 2.35 4.18 7.01 5.67
NF 3.32 2.39 2.12 5.14 7.88 6.31
SDF-GAN 3.86 2.73 2.44 4.77 6.87 5.24
Ours 4.21 3.07 2.77 4.00 5.71 5.36
w/ O1 4.31 3.15 2.90 3.72 5.33 5.27
Table 1: Sharpe Ratio (SR) and Maximal Drawdown (MDD) for Tengency Portfolio (TP), Equal-Weighted (EW), and Value-Weighted (VW) long-short portfolio built based on the baselines and our method with the default GPT-4o-0806 or O1-Preview. We bolded the best results and underlined the second bests.
We begin by testing the Sharpe ratio for portfolios built on the predicted returns of individual assets. The Sharpe ratio (SR) Sharpe (1998) quantifies the risk-adjusted performance of a portfolio as $S_{p}=\frac{\bar{r}_{p}-\bar{r}_{f}}{\sigma(r_{p})}$, where $r_{f}$ stands for the risk-free return, $r_{p}$ represents the portfolio return and $\sigma$ indicates the standard deviation. Furthermore, we evaluate the maximum drawdown, which is the largest decrease in the total value of the portfolio up to time $T$, expressed as $MDD(T)=\max_{\tau\in(0,T)}[\max_{t\in(0,\tau)}X(t)-X(\tau)]$. Here, $X(\tau)$ is the highest value, and $X(t)$ is the lowest value of the portfolio within the time interval $(0,\tau)$.
@@
The experiment results are presented in Table 1. Our approach achieved the highest SR across all three portfolios, with SR improvements of 9.0%, 11.2%, and 11.7% respectively over the best baseline methods (SDF-GAN for TP, IPCA for EW, and VW), averaging a 10.6% increase. Additionally, it secured the best or second-best MDD in TP and EW compared to the leading baseline IPCA with gains of 4.3% and 9.7%. In VW, the MDD underperforms the top baseline CA by only 2.3%. However, substituting GPT-4o-0806 with O1-Preview Jaech et al. (2024), which trained with reasoning enhancements, resulted in SR improvements of 9.1%, 14.1%, and 12.9% across the three portfolios, and improved MDD levels to gains of 11.0%, 15.7%, and -0.6% relative to the best baselines.
@@ 4.3 Asset Pricing Error
[…]
NNAP 0.97 3.02 0.67 6.97
IPCA 0.84 2.65 0.60 6.68
CA 0.89 2.62 0.58 6.72
NF 0.93 2.88 0.63 7.56
SDF-GAN 0.80 2.55 0.59 6.67
Ours 0.72 2.47 0.55 6.48
w/ O1 0.69 2.41 0.55 6.32
Table 2: Asset pricing errors for anomaly portfolios with the baselines and our method with GPT-4o-0806 and O1-Preview. We bolded the best results and underlined the second bests.
@@
Table 2 displays the results. Our method secured either the top or second-best performance among all benchmarks. It demonstrates a 10.0% and 13.8% reduction in average $|\alpha|$ for GPT-4o-0806 and O1-Preview respectively when compared to SDF-GAN, the leading benchmark, along with a 3.1% and 5.5% increase in t-value. Additionally, there is a 5.2% reduction in the proportion of pricing results with a t-value exceeding 1.96 compared to CA for both GPT-4o-0806 and O1-Preview, as well as a 2.8% and 5.3% enhancement in the GRS test compared to SDF-GAN.
@@ 4.4 Ablation Study
SR (EW) MDD (EW) avg $|\alpha|$ avg $|t(\alpha)|$
NF 2.39 7.88 0.97 3.02
+ Factors 2.33 9.12 1.03 3.16
Naive 2.44 6.13 0.95 2.77
+ RAG 2.59 5.99 0.93 2.71
+ Emb. 2.56 5.82 0.92 2.72
Memory 2.75 7.17 0.87 2.69
+ Factors 2.82 6.77 0.84 2.65
Hybrid 2.88 6.32 0.82 2.49
+ Refine 3.01 5.89 0.77 2.32
+ Notes 2.99 6.68 0.76 2.53
Ours 3.07 5.71 0.72 2.47
Table 3: Ablation study of our method and comparison with NF Bybee et al. (2023). “Naive” directly produces the analysis report given news and only daily embeddings are inputted to the pricing network. “+ RAG” introduces the external memory and retrieves Top-K items when performing analysis for Retrieval-Augmented Generation (RAG). “+ Emb.” introduces the asset embeddings. “Memory” baseline incorporate both “+ RAG” and “+ Emb.” “+ Factors” introduces the daily manual factors into the pricing network in “Memory”. “Hybrid” baseline pretrained the pricing network of “Memory”. “+ Refine” refines the analysis report iteratively in $N$ rounds for “Hybrid”. “+ Notes” introduces the macro economics and micro market notes. “Ours” is our method that combines “+ Refine” and “+ Notes” in “Hybrid”.
@@ 4.4.1 Agent Architecture Design
We analyze our architecture in a reverse manner, beginning with a ”Naive” agent that generates the analysis report directly from the refined news without any supplementary information or iterative analysis, while the pricing network solely uses the daily embeddings as input. We then incrementally add components to develop stronger baselines until arriving at our method. The results are shown in Table 3, and the baseline illustrations are provided in Appendix B.
@@
Owing to the analytical capabilities and feature extraction proficiency of LLMs, the “Naive” baseline enhances the SR by 2.1% with comparable pricing errors to NF. Incorporating external memory with RAG further boosts the SR by 6.1% and decreases the average $|\alpha|$ by 2.1% over “Naive”, highlighting the significance of additional contextual information when interpreting business news. Moreover, asset embedding leads to a 4.9% increase in SR and a 3.2% reduction in average $|\alpha|$ by introducing asset-specific loadings.
By combining both, the “Memory” baseline enhances the SR of “Naive” by 12.7% and decreases the average $|\alpha|$ by 8.4% with a lower $t$ value. Incorporating the manual factors, the SR saw a slight increase of 2.5%, while the average $|\alpha|$ decreased by 3.4%. In comparison, the performance of NF declined after the introduction of manual factors, which is consistent with the findings of Bybee et al. (2023), where the inclusion of Fama-French factors reduces the SR, which may be due to suboptimal interactions between factors and news features.
After pretraining the pricing network with historical factor data, the performance of the “Hybrid” baseline saw a notable enhancement of 4.7% in SR and a 5.7% reduction in the average $|\alpha|$ when compared to the “Memory” baseline. This demonstrates the synergy between manual factors and LLM-generated reports, resulting in a successful non-linear interaction. The improvements from our iterative refinement and long-term notes over the “Hybrid” baseline are 4.5% and 3.8% in SR, 4.1% and 6.1% and 7.3% in average $|\alpha|$, respectively with a lower $t$ value and a similar level of MDD. These enhancements collectively yield 6.6% and 12.2% gains in SR and average $|\alpha|$ respectively in our full method compared to a “Hybrid” baseline, underscoring the effectiveness of our agent architecture design.
@@ 4.4.2 Analysis Depth and Width
Figure 4: The Sharpe ratio of equal-weighted portfolios given different numbers of $K$ and $N$.
We further investigate the depth of the analysis, which is controlled by the number of iterations $N$ to refine the analysis report, and the width, which is determined by $K$, the amount of relevant information to check. The results are shown in Figure 4. We keep one variable constant and test the other. We observe that the agent benefits from more rounds of analysis and a broader range of relevant information overall with a sharp decline in marginal gain after a certain point around $K\times N=15$, likely due to the sufficiency of the provided information. Thus, we test an extreme case where $N=1$ and $K=15$, resulting in the SR dropping to 2.98. This indicates that iterative refinement is necessary, as items retrieved in different rounds of refinement provide diverse information as the query evolves. In contrast, a single retrieval leads to items falling into the same topic, with the value of additional items decreasing rapidly and potentially introducing noise.
4.4.3 Foundation Model Impact
SR (EW) MDD (EW) avg $|\alpha|$ avg $|t(\alpha)|$ Input (\$) Output (\$)
GPT-4 3.03 5.57 0.73 2.56 30 60
GPT-3.5 2.89 5.64 0.75 2.60 1.0 2.0
GPT-4o (Ours) 3.07 5.71 0.72 2.47 2.5 10
O1-preview 3.15 5.33 0.69 2.41 15 60
O1-mini 3.12 5.12 0.72 2.56 1.1 4.4
Llama 3.1 405B 3.01 5.29 0.73 2.59 3.5 3.5
Llama 3.1 70B 2.94 5.91 0.75 2.69 0.9 0.9
Qwen 2 72B 2.91 6.89 0.77 2.67 0.9 0.9
Table 4: Ablation study of foundation models.
Table 4 shows the analysis of how selecting different foundation models affects outcomes. We begin by contrasting widely used OpenAI models, including GPT-4o-0806 employed in our study, as well as older models such as GPT-3.5-Turbo-1106 and GPT-4-0613, to evaluate whether advanced foundational models enhance asset pricing precision. A notable finding is a 4.6% increase in SR when comparing GPT-3.5 to GPT-4. Although GPT-4o shows minimal disparity with GPT-4, it is significantly more cost-effective. Our findings indicate that our method can benefit from the evolution of LLMs, possibly yielding better outcomes with newer models over time. Importantly, the outdated models in this experiment have a knowledge cut-off preceding our dataset, thereby eliminating any information leakage risk, yet they still surpass the baselines; specifically, GPT-3.5 and GPT-4 improve SR by 5.9% and 11.0% compared to the leading SDF-GAN and improve avg. $|\alpha|$ by 6.3% and 8.8% respectively.
Additionally, recent progress in reasoning models, such as O1-preview and O1-mini, demonstrates a 2.6% and 1.6% enhancement in SR compared to standard chat models like GPT-4o. Experiments also include open-source models via together.ai  [^3] https://www.together.ai/, with notable contenders being Llama 3.1 and Qwen 2 across various sizes. Of these, the formidable Llama 3.1 405B rivals GPT-4 at a reduced token cost, and smaller variants around 70B surpass the older GPT-3.5, again offering a cost advantage. These findings underline the promise of high-efficiency local deployment using our method.
@@ 5 Discussion
Our proposed approach presents a promising method to fuse qualitative discretionary investment with quantitative factor-based strategies through the use of LLM agents in asset pricing. Nonetheless, there is still much to investigate regarding additional capabilities of LLM agents that could further enhance asset pricing power. Firstly, internet access and a broader range of information sources, including those available in SocioDojo, may enable the agent to generate more in-depth analyses, as discretionary investment relies on information beyond just news or domain knowledge. Secondly, employing specialized financial LLMs like FinGPT Yang et al. (2023) could further improve the agent’s financial analytical capabilities. Finally, it is crucial to consider multimodal information, such as diagrams and figures, which are frequently presented in financial documents.
@@ 6 Conclusion
In this research, we introduced an LLM agent asset pricing model, a model that combines qualitative analysis from the LLM agent with quantitative factors in asset pricing. Our method surpassed established asset pricing methods in multiple evaluations, including portfolio optimization and asset pricing error. Additionally, we performed an in-depth analysis of each component in our agent design. We posit that our research enhances the understanding of how discretionary investment strategies intersect with quantitative factor-driven models, providing evidence for the effective application of LLMs in asset pricing, and contributing to societal improvements in economic efficiency.
@@
Reproducibility Statement
Comprehensive descriptions of our experimental configuration are presented in Section 4.1, while details pertaining to hyperparameter optimization are provided in Appendix A. The ablation study methodology is elaborated upon in Appendix B. We document all utilized resources, such as the CRSP database and the together.ai platform. Subsequent to publication, our code and data for replicating the study outcomes will be made accessible.